# Light-Weight Diffusion Multiplier and Uncertainty Quantification for Fourier Neural Operators

### v1

@@ Abstract
Operator learning is a powerful paradigm for solving partial differential equations, with Fourier Neural Operators serving as a widely adopted foundation. However, FNOs face significant scalability challenges due to overparameterization and offer no native uncertainty quantification – a key requirement for reliable scientific and engineering applications. Instead, neural operators rely on post hoc UQ methods that ignore geometric inductive biases. In this work, we introduce DINOZAUR: a diffusion-based neural operator parametrization with uncertainty quantification. Inspired by the structure of the heat kernel, DINOZAUR replaces the dense tensor multiplier in FNOs with a dimensionality-independent diffusion multiplier that has a single learnable time parameter per channel, drastically reducing parameter count and memory footprint without compromising predictive performance. By defining priors over those time parameters, we cast DINOZAUR as a Bayesian neural operator to yield spatially correlated outputs and calibrated uncertainty estimates. Our method achieves competitive or superior performance across several PDE benchmarks while providing efficient uncertainty quantification.
@@ 1 Introduction
Partial differential equations (PDEs) are a fundamental mathematical tool that describe a wide range of physical phenomena observed in the real world. Numerous problems in science and engineering require formulating and solving PDEs, often presenting significant computational and methodological challenges to researchers and practitioners. A common strategy for solving PDEs is via conventional numerical methods, such as finite difference, finite element, or finite volume method Larsson and Thomée [2003]. While accurate, they become prohibitively expensive in complex engineering tasks, especially where rapid iteration or real-time response is critical, such as design optimization or control Martins and Ning [2021].
In recent years, there has been a growing trend toward replacing computationally intensive numerical solvers with neural network-based surrogate models trained to approximate PDE solutions. Since discretized PDE domains are commonly represented using point clouds, meshes, or graphs, geometric deep learning Bronstein et al. [2021] has emerged as a promising paradigm for this class of problems. In particular, message passing layers, initially developed for geometric and relational data, have proven effective due to their structural resemblance to physical convolution operators. These layers are especially suitable for capturing the spatial dependencies inherent in PDE systems, making them well-suited for learning dynamics over irregular or non-Euclidean domains Pfaff et al. [2020], Brandstetter et al. [2022].
However, geometric deep learning methods often lack key properties required for solving PDEs, motivating alternative approaches. A PDE – its parameters, equation structure, and boundary conditions – defines an operator that maps input functions to solution functions. Neural operators (NOs) Lu et al. [2019], Kovachki et al. [2023, 2024] have recently emerged as a leading framework for approximating such solution operators and, more generally, for tackling the operator learning problem implicitly present in PDEs. These models incorporate inductive biases that improve performance and offer discretization invariance and universal approximation Kovachki et al. [2023], making them a strong alternative to geometric deep learning. A prominent example is the Fourier Neural Operator (FNO) Li et al. [2021], which learns transformations in the spectral domain via the Fast Fourier Transform. While FNOs are restrictive in requiring input data on a regular grid, they remain a widely used backbone in neural operator architectures Li et al. [2024]. Despite their success on problems such as Darcy flow and Burgers’ equation, FNOs are known to be overparameterized Tran et al. [2023], Kossaifi et al. [2024]. Their design leads to exponential growth with domain dimensionality, causing overfitting and significant memory overhead.
@@
Moreover, PDE solvers often inform downstream decision-making, so it is imperative for any substitute method to quantify the uncertainty on its output. Uncertainty quantification (UQ) is a common strategy to improve model robustness by estimating confidence in its predictions. It enhances model outputs by accounting for different sources of error, helping guide more informed and reliable downstream use. Uncertainty modeling has long relied on Bayesian inference Mackay [1992], ensembles Dong et al. [2020], and conformal prediction Shafer and Vovk [2008] to produce confidence estimates. While some of these methods have been adapted to deep learning with varying success Gal and Ghahramani [2016], Zhang et al. [2021], Jospin et al. [2022], UQ for GNNs and neural operators Psaros et al. [2023] remains underdeveloped, often relying on post hoc methods ill-suited to their structure.
To address the scalability issues and limited UQ methods in FNOs, we present DINOZAUR – a diffusion-based neural operator parametrization with uncertainty quantification. Our diffusion multiplier is inspired by the heat kernel in the Fourier basis and requires only one learnable parameter per channel, resulting in superior scalability without sacrificing predictive performance. This physically motivated choice of multiplier allows us to place meaningful probabilistic priors on the learnable time parameters, recover the posterior via inexpensive variational inference (VI) Blei et al. [2017] of these parameters, and efficiently sample spatially correlated functions. We summarize our contributions as follows:
@@ 2 Related work
The area of designing learnable approximation for PDE solutions is vast and rapidly growing Wang et al. [2024], Huang et al. [2025]. Given the increasing availability of data, the need to incorporate inductive biases, and the inherently geometric structure of many physical problems, two classes of deep learning models have come to dominate the field: geometric deep learning models and neural operators.
@@
Graph neural networks (GNNs) Scarselli et al. [2008] offer a flexible framework for learning on non-Euclidean domains. Many GNNs implement spatial convolution by performing learnable aggregation of local neighborhoods Niepert et al. [2016], Chamberlain et al. [2021]. Neural PDE solvers increasingly rely on spatial GNN layers to model and predict physical phenomena across various domains, including control Li et al. [2019], deformable material simulation Sanchez-Gonzalez et al. [2020], and spatiotemporal mesh evolution Pfaff et al. [2020]. Specific architectures target building robust autoregressive solvers Brandstetter et al. [2022] and encoding the geometric boundary conditions Mayr et al. [2023]. However, their performance often suffers from sensitivity to discretization, limiting generalization across mesh resolutions. Spectral GNNs define convolutions using the graph Laplacian’s eigen-decomposition Bruna et al. [2014], Kipf and Welling [2017], often with polynomial approximations of spectral filters Defferrard et al. [2016], Levie et al. [2018]. These have seen success in 3D shape analysis, particularly with functional maps Ovsjanikov et al. [2012]. DiffusionNet Sharp et al. [2022] exemplifies this approach by using the spectral form of the heat kernel for a simple and interpretable kernel parametrization. It exhibits strong performance and generalization; we note that such a kernel is applicable in arbitrary domains.
@@
Neural operators are a powerful class of models for learning mappings between infinite-dimensional spaces in a resolution-invariant manner Kovachki et al. [2023], offering an alternative to GNNs. Lu et al. [2019] introduced DeepONet, inspired by operator learning and universal approximation theory, using a branch-trunk architecture to encode inputs and evaluate the solution at arbitrary locations. A parallel line of work introduced the Graph Neural Operator (GNO) architecture Anandkumar et al. [2020], Li et al. [2020], transferring the message passing intuition from GNNs to NOs. A family of neural operators decomposes the input function into a basis defined by one of the popular transforms: Fourier Li et al. [2021], Laplace Cao et al. [2024], Wavelet Tripura and Chakraborty [2023]. In particular, FNOs are popular for their performance and interpretability despite needing a regular grid sampling, and are commonly used in other networks, such as Geo-FNO Li et al. [2023], U-FNO Wen et al. [2022], and Geometry-Informed Neural Operator (GINO) Li et al. [2024], to name a few. However, FNOs face scalability and generalization issues due to exponential parameter growth with domain dimensionality Tran et al. [2023]. Efforts to improve this, such as FFT factorization Tran et al. [2023], low-rank tensor decomposition Kossaifi et al. [2024], and MLPs in the spectral domain Xiao et al. [2024], have reduced growth but still rely on spatial dimensions.
@@
UQ in deep learning has been widely studied, with popular methods including MC Dropout Gal and Ghahramani [2016], Laplace approximation MacKay [1992], ensembles Rahaman et al. [2021], and non-parametric conformal prediction Zhang et al. [2021], which typically require extra computations or careful calibration. In the context of functional UQ – crucial for neural operators – the landscape is similarly limited. Recent approaches include applying a post hoc Laplace approximation to the final layer of the NO Magnani et al. [2022], using conformal methods after training Ma et al. [2024], or learning an uncertainty measure as an auxiliary output Bülte et al. [2025]. However, none of these strategies model uncertainty within the NO; instead, they adopt generic methods for neural networks and neglect the inductive biases of neural operators.
@@ 3.1 Problem setting
where $f(x)$ is a fixed forcing function and $g(x)$ is a boundary condition. This formulation is general enough to include varying geometric boundaries by including the signed distance function of a geometry in functions from $\mathcal{A}$ and defining a solution as an extension to the whole domain Stein [1970], Li et al. [2024].
@@
where $\kappa_{\gamma_{i}}$ is a kernel, matrix $W^{\mathrm{skip}}\in\mathbb{R}^{d_{c}\times d_{c}}$, $b_{i}\in\mathbf{R}^{d_{c}}$ is a bias and $d_{c}$ is the number of channels. FNO restricts domain to be periodic torus $\Omega=\mathbb{T}^{d}$, kernels to be stationary and represents the integral transform $\mathcal{I}$ in the following way:
@@ 3.2 DINOZAUR architecture
Motivated by the success of DiffusionNet Sharp et al. [2022] and its resemblance to FNO, we choose to restrict the form of $\kappa_{\gamma}$ to that of the heat kernel for our architecture. Consider the heat equation on the torus $\Omega=\mathbb{T}^{d}$:
@@
with a solution Strauss [2007] given by:
@@
$v_{i+1}(x)=\sigma\left(W^{\mathrm{skip}}_{i}v_{i}(x)+b_{i}+W_{i}^{\mathrm{mix}}\begin{bmatrix}\mathcal{I}_{\tau_{i}}[v_{i}](x)\\ \big{(}\mathcal{G}_{\theta_{i}}\circ\mathcal{I}_{\tau_{i}}\big{)}[v_{i}](x)\end{bmatrix}\right).$ (8)
@@ 3.3 Bayesian inference for diffusion multiplier
where $f(\cdot)$ denotes functions evaluated at any finite number of points in the domain $\Omega$, $\mathbf{N}$ denotes a normal distribution, and $\mu_{\text{prior}},\sigma_{\text{prior}}$ are hyperparameters of the prior distribution. We collect all time variables from each channel $c$ and block $i$ for notational convenience, $\tau=\left\{(\tau^{c}_{i})_{c=1}^{d_{c}}\right\}_{i=1}^{M}$. See Figure 2(a) for the corresponding graphical model. We also collect training data in a set of discrete observations, $\mathcal{D}=\left\{\{a_{nk}\}_{k=1}^{K_{n}},\{u_{nj}\}_{j=1}^{J_{n}}\right\}_{n=1}^{N}$, with possibly different sets of evaluation points $\{x_{k}\}$, $\{y_{j}\}$.
@@
We then seek to find $\phi$ that minimizes the Kullback-Leibler divergence (KL) from the $q(\tau;\phi)$ to the true posterior. As this objective is still intractable Murphy [2022], we instead maximize a lower bound (ELBO):
@@
which allows to learn jointly the network parameters $(\theta,\sigma)$ and variational parameters $\phi$. During inference, we replace the true posterior $p_{\theta,\sigma}$ with its approximation $q(\tau;\phi)$ in (10).
@@ 4 Experiments
Table 1: Comparison of DINOZAUR against the baselines in deterministic setting, reporting the $\mathrm{RL}_{2}$ values on test split. The best result is bold, second best is $\underline{\text{underlined}}$
Model # params $\mathrm{RL}_{2}\downarrow$ params $\mathrm{RL}_{2}\downarrow$
(×$(\times 10^{3})$ (×$(\times 10^{-2})$ (×$(\times 10^{3})$ (×$(\times 10^{-2})$
@@
FNO 361.4 $\underline{0.64}$ 12,650.8 $\underline{15.96}$
FNOg 371.9 0.86 12,692.2 17.24
TFNO 17.7 4.44 69.1 21.39
TFNOg 28.2 2.28 110.6 19.40
@@
DINOZAUR (Ours) 28.0 0.39 109.7 15.86
@@
GINO 49,608.6 8.30 49,612.9 8.51
GINOg 49,650.1 7.03 49,645.3 $\underline{8.07}$
TGINO 149.8 7.42 154.0 8.13
TGINOg 191.3 8.09 195.5 8.86
@@
DINOZAUR (Ours) 180.4 $\underline{7.08}$ 184.6 7.96
We evaluate the performance of the DINOZAUR architecture on several standard benchmarks. Section 4.1 presents results for the deterministic variant of DINOZAUR. In Section 4.2, we enable the Bayesian version and assess its UQ capabilities, comparing it to established UQ methods for neural operators. Ablation studies and computational performance comparisons are detailed in Section 4.3. All models are implemented in PyTorch Paszke et al. [2017] and trained on a single NVIDIA A100 GPU for 150 epochs with AdamW Loshchilov and Hutter [2019] and One-Cycle scheduler Smith and Topin [2019] or AdamW Schedule-Free Defazio et al. [2024] with learning rate (in most cases) set to $10^{-3}$. See the Appendix B for details.
@@
We evaluate our method on two regularly sampled datasets from Li et al. [2021] and two datasets with irregular domains used in Li et al. [2024]. For the 2D Darcy flow, we learn a mapping from a scalar permeability field of resolution $241^{2}$ to a scalar pressure field, applying standard scaling to both. The 2D Navier-Stokes (NS) dataset contains $64^{2}$ spatial points and 20 time steps of simulations with viscosity $\nu=10^{-5}$ on a uniform grid. We learn to map the first 10 time steps of vorticity to the last 10, without applying data scaling. For further dataset and PDE details, we refer to Takamoto et al. [2022]. The irregular-domain benchmarks include two datasets for predicting surface pressure on 3D meshes: ShapeNet car Umetani and Bickel [2018] and Ahmed bodies Ahmed et al. [1984]. Following the protocol of Li et al. [2024], we evaluate the signed distance function (SDF) of each mesh on a $64^{3}$ uniform grid and apply standard scaling to the outputs during training. We learn a map from SDF values to scalar pressure on mesh vertices, with inlet velocity added to input in the Ahmed dataset. A summary of the benchmark datasets is provided in the Appendix A.

### v2

@@ Abstract
Operator learning is a powerful paradigm for solving partial differential equations, with Fourier Neural Operators serving as a widely adopted foundation. However, FNOs face significant scalability challenges due to overparameterization and offer no native uncertainty quantification – a key requirement for reliable scientific and engineering applications. Instead, neural operators rely on post hoc UQ methods that ignore geometric inductive biases. In this work, we introduce DINOZAUR: a diffusion-based neural operator parametrization with uncertainty quantification. Inspired by the structure of the heat kernel, DINOZAUR replaces the dense tensor multiplier in FNOs with a dimensionality-independent diffusion multiplier that has a single learnable time parameter per channel, drastically reducing parameter count and memory footprint without compromising predictive performance. By defining priors over those time parameters, we cast DINOZAUR as a Bayesian neural operator to yield spatially correlated outputs and calibrated uncertainty estimates. Our method achieves competitive or superior performance across several PDE benchmarks while providing efficient uncertainty quantification. The code is available at https://github.com/PhysicsXLtd/DINOZAUR.
@@ 1 Introduction
Partial differential equations (PDEs) are a fundamental mathematical tool that describe a wide range of physical phenomena observed in the real world. Numerous problems in science and engineering require formulating and solving PDEs, often presenting significant computational and methodological challenges to researchers and practitioners. A common strategy for solving PDEs is via conventional numerical methods, such as finite difference, finite element, or finite volume method [1]. While accurate, they become prohibitively expensive in complex engineering tasks, especially where rapid iteration or real-time response is critical, such as design optimization or control [2].
In recent years, there has been a growing trend toward replacing computationally intensive numerical solvers with neural network-based surrogate models trained to approximate PDE solutions. Since discretized PDE domains are commonly represented using point clouds, meshes, or graphs, geometric deep learning [3] has emerged as a promising paradigm for this class of problems. In particular, message passing layers, initially developed for geometric and relational data, have proven effective due to their structural resemblance to physical convolution operators. These layers are especially suitable for capturing the spatial dependencies inherent in PDE systems, making them well-suited for learning dynamics over irregular or non-Euclidean domains [4, 5].
However, geometric deep learning methods often lack key properties required for solving PDEs, motivating alternative approaches. A PDE – its parameters, equation structure, and boundary conditions – defines an operator that maps input functions to solution functions. Neural operators (NOs) [6, 7, 8] have recently emerged as a leading framework for approximating such solution operators and, more generally, for tackling the operator learning problem implicitly present in PDEs. These models incorporate inductive biases that improve performance and offer discretization invariance and universal approximation [7], making them a strong alternative to geometric deep learning. A prominent example is the Fourier Neural Operator (FNO) [9], which learns transformations in the spectral domain via the Fast Fourier Transform. While FNOs are restrictive in requiring input data on a regular grid, they remain a widely used backbone in neural operator architectures [10]. Despite their success on problems such as Darcy flow and Burgers’ equation, FNOs are known to be overparameterized [11, 12]. Their design leads to exponential growth with domain dimensionality, causing overfitting and significant memory overhead.
@@
Moreover, PDE solvers often inform downstream decision-making, so it is imperative for any substitute method to quantify the uncertainty on its output. Uncertainty quantification (UQ) is a common strategy to improve model robustness by estimating confidence in its predictions. It enhances model outputs by accounting for different sources of error, helping guide more informed and reliable downstream use. Uncertainty modeling has long relied on Bayesian inference [13], ensembles [14], and conformal prediction [15] to produce confidence estimates. While some of these methods have been adapted to deep learning with varying success [16, 17, 18], UQ for GNNs and neural operators [19] remains underdeveloped, often relying on post hoc methods ill-suited to their structure.
To address the scalability issues and limited UQ methods in FNOs, we present DINOZAUR – a diffusion-based neural operator parametrization with uncertainty quantification. Our diffusion multiplier is inspired by the heat kernel in the Fourier basis and requires only one learnable parameter per channel, resulting in superior scalability without sacrificing predictive performance. This physically motivated choice of multiplier allows us to place meaningful probabilistic priors on the learnable time parameters, recover the posterior via inexpensive variational inference (VI) [20] of these parameters, and efficiently sample spatially correlated functions. We summarize our contributions as follows:
@@ 2 Related work
The area of designing learnable approximation for PDE solutions is vast and rapidly growing [21, 22]. Given the increasing availability of data, the need to incorporate inductive biases, and the inherently geometric structure of many physical problems, two classes of deep learning models have come to dominate the field: geometric deep learning models and neural operators.
@@
Graph neural networks (GNNs) [23] offer a flexible framework for learning on non-Euclidean domains. Many GNNs implement spatial convolution by performing learnable aggregation of local neighborhoods [24, 25]. Neural PDE solvers increasingly rely on spatial GNN layers to model and predict physical phenomena across various domains, including control [26], deformable material simulation [27], and spatiotemporal mesh evolution [4]. Specific architectures target building robust autoregressive solvers [5] and encoding the geometric boundary conditions [28]. However, their performance often suffers from sensitivity to discretization, limiting generalization across mesh resolutions. Spectral GNNs define convolutions using the graph Laplacian’s eigen-decomposition [29, 30], often with polynomial approximations of spectral filters [31, 32]. These have seen success in 3D shape analysis, particularly with functional maps [33]. DiffusionNet [34] exemplifies this approach by using the spectral form of the heat kernel for a simple and interpretable kernel parametrization. It exhibits strong performance and generalization; we note that such a kernel is applicable in arbitrary domains.
@@
Neural operators are a powerful class of models for learning mappings between infinite-dimensional spaces in a resolution-invariant manner [7], offering an alternative to GNNs. Lu et al. [6] introduced DeepONet, inspired by operator learning and universal approximation theory, using a branch-trunk architecture to encode inputs and evaluate the solution at arbitrary locations. A parallel line of work introduced the Graph Neural Operator (GNO) architecture [35, 36], transferring the message passing intuition from GNNs to NOs. A family of neural operators decomposes the input function into a basis defined by one of the popular transforms: Fourier [9], Laplace [37], Wavelet [38]. In particular, FNOs are popular for their performance and interpretability despite needing a regular grid sampling, and are commonly used in other networks, such as Geo-FNO [39], U-FNO [40], and Geometry-Informed Neural Operator (GINO) [10], to name a few. However, FNOs face scalability and generalization issues due to exponential parameter growth with domain dimensionality [11]. Efforts to improve this, such as FFT factorization [11], low-rank tensor decomposition [12], and MLPs in the spectral domain [41], have reduced growth but still rely on spatial dimensions.
@@
UQ in deep learning has been widely studied, with popular methods including MC Dropout [16], Laplace approximation [42], ensembles [43], and non-parametric conformal prediction [17], which typically require extra computations or careful calibration. In the context of functional UQ – crucial for neural operators – the landscape is further limited. Recent approaches include applying a post hoc Laplace approximation to the final layer of the NO [44], using conformal methods after training [45], or learning an uncertainty measure as an auxiliary output [46]. Another family of approaches leverages diffusion-based generative models for uncertainty quantification in physics surrogates [47, 48, 49, 50]. While these methods use diffusion models to capture statistics of high-dimensional dynamical systems and generate realistic samples, they require solving a differential equation for reverse diffusion numerically through multiple denoising steps, which is computationally intensive compared to deterministic forward passes in neural operators. Despite these various approaches, none of them model uncertainty within the NO itself; instead, they adopt generic methods for neural networks and neglect the inductive biases of neural operators.
@@ 3.1 Problem setting
where $f(x)$ is a fixed forcing function and $g(x)$ is a boundary condition. This formulation is general enough to include varying geometric boundaries by including the signed distance function of a geometry in functions from $\mathcal{A}$ and defining a solution as an extension to the whole domain [51, 10].
@@
where $\kappa_{\gamma_{i}}$ is a kernel, matrix $W^{\mathrm{skip}}\in\mathbb{R}^{d_{c}\times d_{c}}$, $b_{i}\in\mathbb{R}^{d_{c}}$ is a bias and $d_{c}$ is the number of channels. FNO restricts domain to be periodic torus $\Omega=\mathbb{T}^{d}$, kernels to be stationary and represents the integral transform $\mathcal{I}$ in the following way:
@@ 3.2 DINOZAUR architecture
Motivated by the success of DiffusionNet [34] and its resemblance to FNO, we choose to restrict the form of $\kappa_{\gamma}$ to that of the heat kernel for our architecture. Consider the heat equation on the torus $\Omega=\mathbb{T}^{d}$:
@@
with a solution [52] given by:
@@
[…]
@@ 3.3 Bayesian inference for diffusion multiplier
where $f(\cdot)$ denotes functions evaluated at any finite number of points in the domain $\Omega$, $\mathbf{N}$ denotes a normal distribution, and $\mu_{\text{prior}},\sigma_{\text{prior}}$ are hyperparameters of the prior distribution. We collect all time variables from each channel $c$ and block $i$ for notational convenience, $\tau=\left\{(\tau^{c}_{i})_{c=1}^{d_{c}}\right\}_{i=1}^{M}$. See Figure 2(a) for the corresponding graphical model. We also collect training data in a set of discrete observations, $\mathcal{D}=\left\{\{a_{nk}\}_{k=1}^{K_{n}},\{u_{nj}\}_{j=1}^{J_{n}}\right\}_{n=1}^{N}$, with possibly different sets of function evaluation points at the input and output.
@@
We then seek to find $\phi$ that minimizes the Kullback-Leibler divergence (KL) from $q(\tau;\phi)$ to the true posterior. As this objective is still intractable [53], we instead maximize a lower bound (ELBO):
@@
which allows to jointly learn the network parameters $(\theta,\sigma)$ and variational parameters $\phi$. To predict for a new instance, we replace the true posterior $p_{\theta,\sigma}$ with its approximation $q(\tau;\phi)$ in (10).
@@ 4 Experiments
Table 1: Comparison of DINOZAUR against the baselines in deterministic setting, reporting the $\mathrm{RL}_{2}$ values on test split. Mean and standard deviation values are computed from three trained models. The best result is bold, second best is $\underline{\text{underlined}}$
Model # params $\mathrm{RL}_{2}\downarrow$ # params $\mathrm{RL}_{2}\downarrow$
$(\times 10^{3})$ $(\times 10^{-2})$ $(\times 10^{3})$ $(\times 10^{-2})$
@@
FNO 361.4 $\underline{0.90\pm 0.03}$ 12,650.8 $\underline{15.90\pm 0.20}$
FNOg 371.9 $0.91\pm 0.07$ 12,692.2 $17.08\pm 0.12$
TFNO 17.7 $8.04\pm 0.86$ 69.1 $21.48\pm 0.27$
TFNOg 28.2 $4.16\pm 0.61$ 110.6 $19.47\pm 0.58$
@@
DINOZAUR (Ours) 28.0 $\mathbf{0.75\pm 0.10}$ 109.7 $\mathbf{15.68\pm 0.21}$
@@
GINO 49,608.6 $8.03\pm 0.39$ 49,612.9 $8.57\pm 0.13$
GINOg 49,650.1 $7.76\pm 0.64$ 49,645.3 $8.25\pm 0.16$
TGINO 149.8 $\underline{7.27\pm 0.16}$ 154.0 $\underline{8.19\pm 0.49}$
TGINOg 191.3 $8.30\pm 0.38$ 195.5 $8.23\pm 0.64$
@@
DINOZAUR (Ours) 180.4 $\mathbf{7.04\pm 0.06}$ 184.6 $\mathbf{7.99\pm 0.07}$
We evaluate the performance of the DINOZAUR architecture on several standard benchmarks. Section 4.1 presents results for the deterministic variant of DINOZAUR. In Section 4.2, we enable the Bayesian version and assess its UQ capabilities, comparing it to established UQ methods for neural operators. Ablation studies and computational performance comparisons are detailed in Section 4.3. All models are implemented in PyTorch [54] and trained on a single NVIDIA A100 GPU for 150 epochs with AdamW [55] and One-Cycle scheduler [56] or AdamW Schedule-Free [57] with learning rate (in most cases) set to $10^{-3}$. See the Appendix B for details. All experiments were repeated three times with random initialization of trainable parameters. Reported results include the mean and standard deviation over all runs.
@@
We evaluate our method on two regularly sampled datasets from [9] and two datasets with irregular domains used in [10]. For the 2D Darcy flow, we learn a mapping from a scalar permeability field of resolution $241^{2}$ to a scalar pressure field, applying standard scaling to both. The 2D Navier-Stokes (NS) dataset contains $64^{2}$ spatial points and 20 time steps of simulations with viscosity $\nu=10^{-5}$ on a uniform grid. We learn to map the first 10 time steps of vorticity to the last 10, without applying data scaling. For further dataset and PDE details, we refer to [58]. The irregular-domain benchmarks include two datasets for predicting surface pressure on 3D meshes: ShapeNet car [59] and Ahmed bodies [60]. Following the protocol of Li et al. [10], we evaluate the signed distance function (SDF) of each mesh on a $64^{3}$ uniform grid and apply standard scaling to the outputs during training. We learn a map from SDF values to scalar pressure on mesh vertices, with inlet velocity added to input in the Ahmed dataset. A summary of the benchmark datasets is provided in the Appendix A.